OpenStudy (anonymous):

what is torque qualitatively and quantitatively?please help me .....i just need it asap.....thanks:)

OpenStudy (aravindg):

what did u mean actually by the qn?

OpenStudy (anonymous):

Torque is a rotational force. for example, here two equal forces are acting in opposite directions, and hence net force on the system is zero. but the object will with no translatory motion. having said that, torque = force X perpendicular distance at a point. = F X r (cross product)

OpenStudy (anonymous):

\[\tau=r X F\] All the quantities above are vectors . It is nothing but the twisting action of a forces applied in opposite directions. At any point on the circumference of a circle, the r and F are perpendicular because the r vector is along the radius of the circular trajectory and the F vector is tangential to the circular trajectory. so \[\tau=\left| r \right|\left| F \right|\sin \theta\] here theta=90 degrees sine 90=1 So τ=|r||F| this is about the magnitude of the torque.